OpenStudy (anonymous):

The slope of a line is 2/3 What is the slope of a line that is parallel to it?

OpenStudy (tukitw):

The gradients or slopes of two parallel lines are the same.

OpenStudy (anonymous):

true...parallel lines have the same slope

OpenStudy (anonymous):

A. 3/2 B. -2/3 C.-3/2 D.2/3

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

the same slope

OpenStudy (anonymous):

help in another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The graph of which of the following will be parallel to the graph of 4x – 3y = –12?

OpenStudy (anonymous):

do you know what the slope of that equation is ?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

Put the equation in y = mx + b form, and m is your slope. Can you do that ?

OpenStudy (anonymous):

idk how to put it in y=mx + b

OpenStudy (anonymous):

4x - 3y = -12 we want the y term on one side and everything else on the other side. so we will add 3y to both sides 4x - 3y + 3y = -12 + 3y -- combine like terms 4x = -12 + 3y -- now we will add 12 to both sides 4x + 12 = -12 + 12 + 3y -- combine like terms 4x + 12 = 3y -- now we will divide both sides by 3 4/3x + 12/3 = y.......y = 4/3x + 4 as you can see, the number in the m position is 4/3. A parallel line will have the same slope. So the parallel line will have a slope of 4/3

OpenStudy (anonymous):

do you understand ? I tried to leave all the steps in so you could follow

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

The graph of which of the following will be parallel to the graph of y = 2x – 4? A. 2x + y = 4 B. 4x – 2y = –12 C. y = –4x + 8 D. y = 3x – 4

OpenStudy (anonymous):

y = 2x - 4 In y = mx + b form, the slope is in the m position. What is the slope ?

OpenStudy (anonymous):

you can do this...what number is in the m position ?

OpenStudy (anonymous):

skullpatrol...can you explain this any easier for Roberto ?

OpenStudy (anonymous):

I am sorry Roberto, but I have to get ready to go somewhere. I wish I could stay and help, but I have to go somewhere. Your slope is 2. You will need to find the other equation that has 2 as the slope. To do this, put the equations in y = mx + b form, and whichever one has 2 as the slope is the parallel line. Sorry...I have to go now :(

OpenStudy (anonymous):

thanks